# SkyDiffusion: Street-to-Satellite Image Synthesis with Diffusion Models and BEV Paradigm

### v2

@@ Methods
Following previous work (Ke et al. 2022, 2023), we apply different LUT filters to the satellite images to simulate the results under various lighting conditions for training the light manipulation module. During training, we use L2 consistency loss $\mathcal{L}_{\text{con}}$ to maintain image content consistency and L1 reconstruction loss $\mathcal{L}_{\text{rec}}$ to learn color style transfer.
@@
X-Seq 161.16 0.084 11.97 0.706 190.12 0.042 12.41 0.661 - - - -
CDTE 122.84 0.143 10.01 0.694 160.81 0.091 12.59 0.663 - - - -
SelectionGAN 116.57 0.129 12.38 0.742 100.21 0.116 12.04 0.684 149.53 0.127 11.77 0.778
Aerial Diff 136.18 0.103 10.06 0.855 127.29 0.108 10.24 0.878 123.16 0.141 11.49 0.831
HawkI 90.98 0.114 11.09 0.783 135.75 0.112 10.97 0.764 123.47 0.124 9.67 0.752
CUT 72.83 0.121 12.09 0.687 62.22 0.102 13.11 0.664 69.42 0.169 13.66 0.665
img2img-turbo 77.95 0.127 12.64 0.685 73.24 0.103 12.19 0.679 80.10 0.135 11.28 0.690
Instr-p2p 38.01 0.138 12.29 0.697 49.62 0.116 12.51 0.682 50.30 0.163 10.62 0.689
Ours 29.18 0.168 14.58 0.635 35.51 0.123 13.58 0.657 45.29 0.186 11.69 0.661
@@
Implementation Details. We implement SkyDiffusion using the pre-trained Stable Diffusion v1.5 model (Rombach et al. 2022), with an unlocked diffusion decoder and a classifier-free guidance scale of 9.0 (Ho and Salimans 2022). Final inference sampling uses 50 steps with the DDIM strategy (Song, Meng, and Ermon 2021). Training is on eight NVIDIA A100 GPUs with a batch size of 128 for 100 epochs. For light manipulation, the reference satellite image is randomly selected from normally toned images in the training set, and the DNCM mapping matrix parameter $k$ is set to 32 (details in Supplementary Material).
Experiments
In this section, we introduce the three datasets utilized and the experimental setup. We then conduct both qualitative and quantitative comparisons of SkyDiffusion with state-of-the-art cross-view synthesis methods. Finally, we perform ablation studies to evaluate the effectiveness of each module.
Datasets
CVUSA & CVACT (Suburban dataset). CVUSA (Workman, Souvenir, and Jacobs 2015) and CVACT (Liu and Li 2019) are two widely used cross-view datasets. They contain centrally aligned satellite and street-view images from regions in the United States and Canberra, Australia, respectively, with a one-to-one mapping relationship. The data primarily consists of suburban scenes characterized by relatively low-rise buildings and open views. The CVUSA dataset’s training and testing sets are randomly split, while the CVACT dataset is split by region.
VIGOR-Chicago (Urban dataset). VIGOR-Chicago (Zhu, Yang, and Chen 2021) is a subset of the VIGOR dataset, containing cross-view image pairs collected from Chicago urban area. In this dataset, the street panorama and satellite images are not centrally aligned, and multiple street-view images cover the same satellite image area, with overlapping regions between different satellite images, forming a multi-to-one mapping relationship. All input street images are set to 512 × 1024 pixels, with the center column aligned to the north. Unlike previous methods that synthesize low-resolution satellite images (64 $\times$ 64 or 256 $\times$ 256 pixels), this experiment aims to meet practical needs by generating higher-resolution satellite images (512 $\times$ 512 pixels). In the Curved-BEV transformation, $\lambda$ is set to 3 and $H$ to 1.5 for CVUSA and CVACT datasets. For the VIGOR-Chicago dataset, $\lambda$ is set to 10, with $H$ unchanged. The Curved-BEV transformation results in street images that match the size of the satellite images.
Experimental Setup
We compared our method on three datasets with several state-of-the-art cross-view synthesis methods, including the commonly used cross-view satellite synthesis methods X-Seq (Regmi and Borji 2018), SelectionGAN (without semantic masks) (Tang et al. 2019), and CDTE (Toker et al. 2021), image translation methods such as the GAN-based CUT (Park et al. 2020) and diffusion-based Instruct pix2pix (Instr-p2p) (Brooks, Holynski, and Efros 2023) and img2img-turbo (Parmar et al. 2024), as well as novel view synthesis methods like Aerial Diff (Kothandaraman et al. 2023a) and HawkI (Kothandaraman et al. 2023b).
Except for the fine-tuned diffusion model, which requires brief textual descriptions, all other models use the same street-view input for a fair comparison.
Following previous work (Lu et al. 2020; Toker et al. 2021; Regmi and Borji 2018), we employ commonly used metrics such as SSIM, PSNR and LPIPS to evaluate the content consistency of the synthesized images, and use FID (Heusel et al. 2017) to assess the realism of images.
Comparison with State-of-the-Art Methods
@@
On the suburban CVUSA and CVACT datasets, our SkyDiffusion method achieved the outstanding results, as shown in Table 1. Compared to state-of-the-art methods, it reduced FID by 25.83% and increased SSIM by 13.89%, demonstrating its superiority in synthesizing realistic and consistent satellite images. The visual results in Fig. 4 show that GAN-based cross-view methods generate excessive artifacts and blurriness. While diffusion-based methods such as Aerial Diff and Instr-p2p produce highly realistic aerial views, they lack content correlation with the street-view images. In contrast, our method benefits from BEV perspective transformation, enabling effective capture of cross-view associative information. SkyDiffusion generates realistic images with consistent content layout, including road direction, tree placement, and building distribution. Additional visual results can be found in the supplementary materials.
@@
In the urban VIGOR-Chicago dataset, SkyDiffusion reduced FID by 9.96% and improved SSIM by 14.11% compared to the state-of-the-art method, as shown in Table 1. For fair comparison, the VIGOR results for SkyDiffusion use only a single street-view image. The results using Multi-to-One BEV generation with multiple street-view images can be found in the ablation study. The visual results in Fig. 4 show that our method maintains consistent associative information such as crosswalks and lanes. While the non-associative information on building rooftops cannot be consistently generated, it still produces reasonable and realistic results.
Ablation Study
@@
CVUSA
Baseline 32.45 0.149 12.63 0.650
Ours (C-BEV) 29.18 0.168 14.58 0.635
CVACT
Baseline 62.21 0.115 11.95 0.682
Ours (C-BEV) 35.51 0.123 13.58 0.657
VIGOR-Chicago
Baseline 53.27 0.170 10.38 0.666
Ours (C-BEV) 45.29 0.186 11.69 0.661
Ours (C-BEV Multi) 31.90 0.205 11.15 0.651
Table 2: Ablation study of the Curved-BEV module. "Baseline" represents directly using street-view image,"C-BEV" denotes using Curved-BEV transformation, and "Multi" stands for Multi-to-One strategy.
Light manipulation. The ablation experiments in Table 3 indicate that the Light Manipulation module aligns the lighting conditions of the synthesized images with those of the target domain images, improving SSIM and PSNR metrics. This is because satellites typically capture large areas with relatively uniform overall brightness. Using a small satellite image slice as a reference effectively captures the lighting conditions at the time of capture. Due to the dispersed distribution of the CVUSA data and the more complex lighting conditions, the improvements are not as pronounced as on other datasets. The light processing module maintains the original image content while providing more flexible lighting conditions for the synthesized satellite images. More visualization results can be found in the supplementary material.
CVUSA
Ours (w/o light) 29.02 0.167 14.07 0.643
Ours 29.18 0.168 14.58 0.635
CVACT
Ours (w/o light) 36.48 0.117 12.85 0.665
Ours 35.51 0.123 13.58 0.657
VIGOR-Chicago
Ours (w/o light) 46.77 0.179 10.46 0.666
Ours 45.29 0.186 11.69 0.661
Table 3: Ablation study of the light manipulation module.
Figure 6: Synthesis results for cross-dataset generalization.
CVACT $\rightarrow$ VIGOR-Chicago
Instr-p2p 100.80 0.152 9.14 0.717
Ours (w/o light) 89.76 0.155 9.45 0.739
Ours 63.04 0.162 11.24 0.723
VIGOR-Chicago $\rightarrow$ CVACT
Instr-p2p 121.29 0.113 8.90 0.759
Ours (w/o light) 107.58 0.116 8.76 0.751
Ours 100.58 0.121 11.36 0.714
Table 4: Cross-dataset generalization assessment.
Cross-Dataset Generalization
We trained the model on CVACT and tested it on the VIGOR-Chicago, and vice versa, to evaluate cross-dataset generation capability, as shown in Table 4. Compared to Instruct pix2pix, our method (w/o light) demonstrates superior performance across metrics. Figure 6 shows the synthesis results for cross-dataset generalization. Our method effectively preserves scene content such as road directions and intersections. However, due to the distinct city styles of the CVACT and VIGOR datasets and the fact that the model was trained on a single dataset, there are noticeable differences in elements such as buildings and lighting conditions. The aim of this task is not to generate identical satellite images but to create scenes with consistent content layouts while exhibiting distinctly different styles. Additionally, we can flexibly use the light manipulation module to obtain light information from a random image in the target dataset’s training set, better matching the target domain’s lighting conditions.
Conclusion
In this study, we introduce SkyDiffusion, a novel approach specifically designed for street images to satellite images cross-view synthesis. SkyDiffusion operates solely with street images as input, utilizing a BEV Paradigm and diffusion models to generate satellite images. SkyDiffusion achieves state-of-the-art performance in both content consistency and image realism on suburban and urban cross-view datasets, demonstrating its superior capabilities in cross-view synthesis tasks. The satellite and street-view domains bridging method, Curved-BEV, within this approach also holds valuable insights for tasks such as cross-view retrieval and geolocation. In the future work, we plan to extend our method to more cities and further refine it for specific downstream applications, such as urban planning and virtual tourism.

### v3

@@ 3.3 Network Training
Implementation Details. We implement SkyDiffusion using the pre-trained Stable Diffusion v1.5 model [25], with an unlocked diffusion decoder and a classifier-free guidance scale of 9.0 [6]. Final inference sampling uses 50 steps with the DDIM strategy [31]. Training is on eight NVIDIA A100 GPUs with a batch size of 128 for 100 epochs. For optional light manipulation, the reference satellite image is randomly selected from normally toned images in the training set, and the DNCM mapping matrix parameter $k$ is set to 32 (details in Supplementary Material).
4 Dataset
We propose Ground2Aerial-3, a multi-task cross-view synthesis dataset designed to explore the performance of cross-view synthesis methods in several novel scenarios. As shown in Figure 3, G2A-3 contains nearly 10k pairs of street-view and aerial images, covering disaster scene aerial image synthesis, historical high-resolution satellite image synthesis, and low-altitude UAV image synthesis. The dataset of each task is randomly split into training and test sets with a ratio of 5:1. The ground street-view images are 1024×512, with true north aligned at the center, and the aerial images are 512×512, aligned with the center of the ground images.
4.1 Disaster Scene Aerial image
Both satellite and aerial images are limited by their acquisition frequency, which imposes certain constraints on rapid response to natural disasters. Synthesizing aerial images based on street-view images captured by ground cameras can facilitate fast disaster assessment and analysis. CVIAN [13] is a cross-view dataset containing street-view and high-resolution satellite imagery in Florida, USA, following Hurricane Ian in 2022. From CVIAN’s street-view and satellite data, we filtered 2.7k image pairs of moderately or severely affected areas and adjusted the north direction of street-view images to the central column, preparing the data for the disaster scene aerial image generation task.
Figure 4: Illustration of the Ground2Aerial-3 dataset.
4.2 Historical High-Resolution Satellite image
Obtaining high-resolution satellite imagery, particularly ultra-high resolution images exceeding 0.3m, has been challenging. It wasn’t until the recent launch of WorldView-3 (2014) and WorldView-4 (2018) that high-resolution coverage at near-global scales became available. Previously, ultra-high resolution images were often limited to small areas captured by aerial photography, with both spatial and temporal coverage constraints. Synthesizing satellite images based on historical street-view images can help fill this gap, providing additional complementary information. We collected historical street-view images from 2007 to 2014 on the Mapillary  [^1] https://www.mapillary.com/ platform and used corresponding aerial imagery from the massGIS platform  [^2] https://www.mass.gov/info-details/massgis-data-layers as ground truth to conduct high-resolution satellite image synthesis tasks.
4.3 Low-Altitude UAV Image
Low-altitude UAV imagery, with ultra-high resolution below 0.05 meters, has potential applications in tasks such as lane data collection for autonomous driving. However, the coverage and cost of UAV imagery are limited, and safety regulations further restrict coverage range and update frequency. Synthesizing low-altitude UAV images from more accessible street-view images presents a potential application scenario. Using virtual MatrixCity [14] data, we employed the UE engine to set up six single-view cameras at ground level, rendering and stitching them into panoramic street-view images. At the same locations, we positioned downward-facing cameras at an altitude of 20 meters to render low-altitude UAV imagery, creating a cross-view dataset for low-altitude UAV image synthesis tasks.
Figure 5: Qualitative comparison of different methods synthesis results on three datasets.
5 Experiments
In this section, we introduce the datasets utilized and the experimental setup. We then conduct both qualitative and quantitative comparisons of SkyDiffusion with state-of-the-art cross-view synthesis methods. Finally, we perform ablation studies to evaluate the effectiveness of each module.
5.1 Experimental Setup
Dataset Details. CVUSA [40] and CVACT [16] are two commonly used cross-view datasets, containing one-to-one, center-aligned aerial and ground street-view images. They mainly covers suburban scenes, with low-rise buildings and open views. The CVACT dataset is regionally segmented. Although CVUSA uses random partitioning, its data distribution is sparse, with no overlap among satellite images.
VIGOR-Chicago [46] is a subset of the VIGOR cross-view dataset, containing the Chicago area. In this dataset, street-view panoramas and satellite images are not center-aligned, with multiple street-view images covering the same satellite image region, forming a multi-to-one mapping relationship. We modified the train-test split to a 4:1 ratio and ensured that the same satellite images do not appear in both the training and testing sets to prevent information leakage. The G2A-3 dataset has been described in Section 4.
All input street images are set to 512 $\times$ 1024 pixels, with the center column aligned to the north. Unlike previous methods that synthesize low-resolution satellite images (64 $\times$ 64), this experiment aims to meet practical needs by generating higher-resolution satellite images (512 $\times$ 512 pixels). In the Curved-BEV transformation, $\lambda$ is set to 2 and $H$ to 1.5 for CVUSA, CVACT and G2A-3 datasets. For the VIGOR-Chicago dataset, when performing a one-to-one mapping, $\lambda$ is set to 10; for a Multi-to-One mapping, $\lambda$ is set to 0, the primary focus is on retaining the street-view central location information. The Curved-BEV transformation generates results that are size-matched with satellite images.
Comparison Methods and Metric. We compared our method with several state-of-the-art cross-view synthesis methods, including the commonly used cross-view satellite synthesis methods X-Seq [23], SelectionGAN (without semantic masks) [33], and CDTE [35], image translation methods such as the GAN-based CUT [20] and diffusion-based Instruct pix2pix (Instr-p2p) [2], ControlNet [43] and img2img-turbo [21], as well as novel view synthesis methods like Aerial Diff [9] and HawkI [10].
Except for the fine-tuned diffusion model, which requires brief textual descriptions, all other models use the same street-view input. For a fair comparison, the results in the table 1 are presented without the use of the Optional Light Manipulation module for Skydiffusion. Following previous work [17, 35, 23], we employ commonly used metrics such as SSIM, PSNR and LPIPS to evaluate the content consistency of the synthesized images, and use FID [5] to assess the realism of images.
@@ 5.1 Experimental Setup
X-Seq [23] 161.16 0.084 11.97 0.706 190.12 0.042 12.41 0.661 - - - -
CDTE [35] 122.84 0.143 10.01 0.694 160.81 0.091 12.59 0.663 - - - -
SelGAN [33] 116.57 0.129 12.38 0.742 100.21 0.116 12.04 0.684 149.53 0.127 11.77 0.778
Aerial Diff [9] 136.18 0.103 10.06 0.855 127.29 0.108 10.24 0.878 123.16 0.141 11.49 0.831
HawkI [10] 90.98 0.114 11.09 0.783 135.75 0.112 10.97 0.764 123.47 0.124 9.67 0.752
CUT [20] 72.83 0.121 12.09 0.687 62.22 0.102 13.11 0.664 69.42 0.169 13.66 0.665
I2I-Turbo [21] 77.95 0.127 12.64 0.685 73.24 0.103 12.19 0.679 80.10 0.135 11.28 0.690
Instr-p2p [2] 38.01 0.138 12.29 0.697 49.62 0.116 12.51 0.682 50.30 0.163 10.62 0.689
ControlNet [43] 32.45 0.149 12.63 0.650 62.21 0.115 11.95 0.682 53.27 0.170 10.38 0.666
Ours 29.18 0.168 14.58 0.635 36.48 0.118 12.85 0.645 45.29 0.186 11.69 0.661
@@
Disaster Scene image synthesis
ControlNet 138.11 0.183 12.95 0.690
Ours 114.64 0.188 13.02 0.694
Historical aerial image synthesis
ControlNet 36.48 0.117 12.85 0.665
Ours 35.51 0.123 13.58 0.657
Low-altitude UAV image synthesis
ControlNet 52.76 0.371 16.55 0.598
Ours 32.11 0.452 16.86 0.390
Table 2: Quantitative comparison of methods on G2A-3 Dataset.
5.2 Comparison with State-of-the-Art Methods
@@
On the suburban CVUSA and CVACT datasets, SkyDiffusion achieved the outstanding results, as shown in Table 1. Compared to state-of-the-art methods, it reduced FID by 25.72% and increased SSIM by 7.68%, demonstrating its superiority in synthesizing realistic and consistent satellite images. The visual results in Fig. 5 show that GAN-based cross-view methods generate excessive artifacts and blurriness. While diffusion-based methods such as Aerial Diff and Instr-p2p produce highly realistic aerial views, they lack content correlation with the street-view images. In contrast, our method benefits from BEV perspective transformation, enabling effective capture of cross-view associative information. SkyDiffusion generates realistic images with consistent content layout, including road direction, tree placement, and building distribution. Additional visual results can be found in the supplementary materials.
Figure 6: Qualitative Visualization Results on the G2A-3 Dataset.
@@
In the urban VIGOR-Chicago dataset, our method reduced FID by 14.98% and improved SSIM by 9.41% compared to the state-of-the-art method, as shown in Table 1. The VIGOR results for SkyDiffusion use only a single street-view image. The results using Multi-to-One BEV generation with multiple street-view images can be found in the ablation study. The visual results in Fig. 5 show that our method maintains consistent associative information such as crosswalks and lanes. While the non-associative information on building rooftops cannot be consistently generated, it still produces reasonable and realistic results.
As shown in Table 2, the tasks on the G2A-3 dataset present certain challenges; however, our method achieves significant performance improvements over the commonly used image-conditioned synthesis method, ControlNet. SkyDiffusion reduces the FID by an average of 19.60% and increases the SSIM by an average of 9.90%. The visual results in Figure 6 demonstrate SkyDiffusion’s ability to synthesize various types of aerial images across different G2A-3 task settings. In Figure 6(a), prominent disaster areas in the street-view image (orange box indicating debris of damaged buildings and red box showing trees downed by the hurricane) are clearly visible in the synthesized satellite image, aiding in subsequent disaster area assessment and analysis. Figure 6(b) shows that SkyDiffusion can synthesize realistic early historical imagery, effectively filling gaps where early high-resolution aerial images are lacking. Figure 6(c) highlights SkyDiffusion’s exceptional performance in low-altitude UAV image synthesis. Due to the lower altitude of UAV captures, there is greater overlap and closer domain similarity between scenes, resulting in superior synthesis quality. SkyDiffusion demonstrates excellent aerial image synthesis capabilities, advancing potential applications in various aerial image synthesis scenarios.
5.3 Ablation Study and More Exploration
@@
C-BEV Light FID ($\downarrow$) SSIM ($\uparrow$) PSNR ($\uparrow$) LPIPS ($\downarrow$)
CVACT 62.21 0.115 11.95 0.682
36.48 0.118 12.85 0.645
35.51 0.123 13.58 0.657
VIGOR 53.27 0.170 10.38 0.666
45.29 0.186 11.69 0.661
(Multi) 37.90 0.193 10.39 0.660
(Multi) 31.90 0.205 11.15 0.651
Table 3: Ablation study of the Curved-BEV and optional Light Manipulation Modules. "Multi" stands for Multi-to-One strategy, and "Light" represents using light manipulation.
Light manipulation. Aerial and satellite images typically capture large areas with relatively uniform overall brightness. Using small aerial image patches as references effectively captures the lighting conditions at the time of capture. We applied an optional Light Manipulation module to align the lighting conditions of the synthesized image with those of the target domain image, as shown in Table 3 The results indicate that Light Manipulation can effectively improve SSIM and PSNR metrics. This module preserves the original image content while providing more flexible lighting conditions for the synthesized satellite image. Additional visualizations can be found in the supplementary material.
6 Conclusion
In this study, we introduce SkyDiffusion, a novel approach specifically designed for street images to satellite images cross-view synthesis. SkyDiffusion operates solely with street images as input, utilizing a BEV Paradigm and diffusion models to generate satellite images. SkyDiffusion achieves state-of-the-art performance in both content consistency and image realism on across multiple cross-view datasets, demonstrating its superior capabilities. Additionally, we introduce a cross-view synthesis dataset, Ground2Aerial-3, featuring aerial image synthesis tasks for multiple new scenes, providing practical value and inspiration for future cross-view synthesis research.